OpenStudy (anonymous):

Evaluate f(x) = x^2 + 1 for x = –3. my answer choices are: A. –9 B. –4 C. –8 D. 4 BUT I GOT f (-3) = 10 ??

OpenStudy (anonymous):

You're right

OpenStudy (anonymous):

(-3)(-3) + 1 = 9 + 1 = 10

OpenStudy (anonymous):

my question answer choices are above 10 is not there :/

OpenStudy (anonymous):

I know, sometimes the question makes a mistake?

OpenStudy (anonymous):

THERE IT GOES

OpenStudy (anonymous):

Oh, see theres a negative sign so its -(-3)(-3) + 1 Then : -(9) + 1

OpenStudy (anonymous):

And what do you get? P:

OpenStudy (anonymous):

-8?!

OpenStudy (anonymous):

Yup